# Evaluating the evaluators: Towards human-aligned metrics for missing markers reconstruction

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Conference; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
Figure 1. Motion capture data from two actors after filling gaps and solving into skeletons. The left and right sides show two different gap-filling methods. This illustration highlights a major issue with commonly used mean squared error metrics for reconstructing missing markers. Although the left prediction looks visually accurate, it shows a higher error than the clearly flawed pose on the right, with broken parts highlighted in light green.
@@ 2.2. Metrics
To incorporate temporal cohesion, we propose using the velocity distance. We suggest two novel metrics based on velocity distance, detailed in Section 4.2.
@@ 3.2. Models considered
This paper focuses on evaluation rather than modeling. We do not present a new model and do not claim to be the new state-of-the-art. Instead, we evaluate two simple models based on commonly used one-dimensional Convolutional Neural Network (CNN) (Kaufmann et al., 2020; Cui and Sun, 2021; Lohit et al., 2021; Xia et al., 2022).
@@
The second model takes a structural approach instead. In contrast to previous structural modeling approaches that make use of graph neural networks (Cui and Sun, 2021; Yin et al., 2021; Lee et al., 2024), we introduce structure mainly during inference by replacing missing markers one body part at a time. Our model consists of four modules, each with the same architecture as the first model. Instead of predicting all markers at once, each module predicts one body part at a time. As shown in Figure 2, it follows the order often used for manual cleanup: hips -¿ torso -¿ head -¿ limbs. Prediction is done iteratively. In each step, we first predict values for the current body part and then overwrite the interpolated or noisy values in $\dot{Y}$ with the corresponding predicted values:
@@ 4.1.1. Data processing
The data was hips-centered by removing an average of the hips’ coordinates from each actor’s data at every frame, so that the hips center for each actor is at the origin (0,0,0) at every frame. The distance between the actors was stored separately at each time frame since this information is lost in the hips-centering step. The input to the model for a sequence of length $T$ is thus of size $[T,2*63*3+3]$.
@@ 4.2.2. Bone distance preservation
(10) $\sqrt{\sum_{t=1}^{T-1}\sum_{d=1}^{D}(\hat{L}_{{t+1},d}-\hat{L}_{td})^{2}}$
Figure 3. Marker set example. Taken from the Vicon website (VICON, [n. d.]).
@@
Figure 4. Screenshot from the data provided in the user study.
We define bone length as the Euclidean distance between the average positions of the two markers placed on different sides of a bone. The following five bones are considered as shown in Figure 3: head (from left to right), left hand (left elbow to wrist), right hand (right elbow to wrist), hips (left to right), left leg (left knee to ankle), right leg (right knee to ankle). For example, for the left-hand bone distance, we first calculate elbow position coordinates as the midpoint between the external and internal elbow markers. We then calculate wrist position as the midpoint between the external and internal wrist markers. Finally, we calculate the left-hand bone distance as the Euclidean distance between the elbow position, and wrist position estimates.
@@ 4.2.3. Velocity distance
To measure temporal coherence, we propose to calculate a velocity distance (VD). This has been previously used in other animation tasks (karras2017audio; holden2020learned), but not for missing marker reconstruction. As in the case of bone distance preservation, there are two ways to define a velocity distance: either with respect to the ground truth or with respect to the previous time frame. We will call the two VD with GT and VD without GT.
@@
(11) $\sqrt{\frac{1}{T}\sum_{t=2}^{T}\lVert v_{t}-\hat{v}_{t}\rVert^{2}}.$
@@
(12) $\sqrt{\frac{1}{T}\sum_{t=2}^{T-1}\lVert\hat{v}_{t+1}-\hat{v}_{t}\rVert^{2}}.$
@@ 6. Correlation between objective metrics and subjective perception
Metric Kendall $\tau$ p-value
RMSE 0.2928 0.0267
VD with GT 0.7345 3e-08
VD 0.3684 0.0052
BDP with GT 0.1538 0.2443
BDP 0.5464 3.358e-05
Table 2. Kendall’s $\tau$ rank correlation coefficients and p-values for different objective metrics wrt the user study results. Bold font indicates statistically significant results (p¡0.01).
@@ Acknowledgements.
Appendix A Hyper-parameters of the models used
Hyper-parameters for the models used are given in Table 3 in the appendix. Velocity coefficient indicates variable $\lambda$ from equation 7 and noise coefficient indicates variable $c$ defining the maximal level of noise in the first paragraph in Section 3.3.
Model # of kernels Kernel size Batch Size Learning Rate Dropout Velocity Coef. Noise Coef.
Vanilla CNN 4096 31 64 1.40E-04 0.0017 4 0.1
Hips Predictor 4096 21 256 5.95E-05 0.209 32 2.56
Torso Predictor 4096 37 64 4.12E-05 0.057 12 1.45
Head Predictor 4096 35 64 6.96E-05 0.15 13 1.65
Limbs Predictor 4096 31 256 2.65E-04 0.0523 12 2.05
Table 3. Main hyper-parameters for the models used in the experiments. Hips Predictor, Torso Predictor, Head Predictor, and Limbs Predictor are the four components of the Hips Outwards model. Note that Vanilla CNN has a very different number for noise coefficient because the data for this model was normalized.

### v3

@@ Abstract.
Animation, Evaluation, Objective Metrics, Motion Capture, Missing Markers Reconstruction, User Studies
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Multimedia; 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755788isbn: 979-8-4007-2035-2/2025/10ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Animationccs: Computing methodologies Machine learning
Figure 1. Motion capture data from two actors after filling gaps and solving into skeletons. The left and right sides show two different gap-filling methods. This illustration highlights a major issue with commonly used mean squared error metrics for reconstructing missing markers. Although the left prediction looks visually accurate, it shows a higher error than the clearly flawed pose on the right, with broken parts highlighted in light green. See the video attachment for an animation example.
@@ 3.2. Models considered
This paper focuses on evaluation rather than modeling. We do not present a new model and do not claim to be the new state-of-the-art. Instead, we evaluate two simple models based on commonly used one-dimensional Convolutional Neural Network (CNN) (Kaufmann et al., 2020; Lohit et al., 2021; Xia et al., 2022).
@@
The second model takes a structural approach instead. In contrast to previous structural modeling approaches that make use of graph neural networks (Cui and Sun, 2021; Yin et al., 2021; Lee et al., 2024), we introduce structure mainly during inference by replacing missing markers one body part at a time. Our model consists of four modules, each with the same architecture as the first model, but each having different hyper-parameters and different network weights. Instead of predicting all markers at once, each module predicts one body part at a time. As shown in Figure 2, it follows the order often used for manual cleanup: hips -¿ torso -¿ head -¿ limbs. Prediction is done iteratively. In each step, we first predict values for the current body part and then overwrite the interpolated or noisy values in $\dot{Y}$ with the corresponding predicted values:
@@ 4.1.1. Data processing
[…]
@@ 4.2.2. Bone distance preservation
(10) $\sqrt{\frac{1}{(T-1)*D}\sum_{t=1}^{T-1}\sum_{d=1}^{D}(\hat{L}_{{t+1},d}-\hat{L}_{td})^{2}}$
@@
We define bone length as the Euclidean distance between the average positions of the two markers placed on different sides of a bone. The following five bones are considered (VICON, [n. d.]): head (from left to right), left hand (left elbow to wrist), right hand (right elbow to wrist), hips (left to right), left leg (left knee to ankle), right leg (right knee to ankle). For example, for the left-hand bone distance, we first calculate elbow position coordinates as the midpoint between the external and internal elbow markers. We then calculate wrist position as the midpoint between the external and internal wrist markers. Finally, we calculate the left-hand bone distance as the Euclidean distance between the left elbow position, and left wrist position estimates.
@@ 4.2.3. Velocity distance
To measure temporal coherence, we propose to calculate a velocity distance (VD). This has been previously used in other animation tasks (Karras et al., 2017; Holden et al., 2020), but not for missing marker reconstruction. As in the case of bone distance preservation, there are two ways to define a velocity distance: either with respect to the ground truth or with respect to the previous time frame. We will call the two VD with GT and VD without GT.
@@
(11) $\sqrt{\frac{1}{(T-1)}\sum_{t=2}^{T}\lVert v_{t}-\hat{v}_{t}\rVert^{2}}.$
@@
(12) $\sqrt{\frac{1}{(T-2)}\sum_{t=2}^{T-1}\lVert\hat{v}_{t+1}-\hat{v}_{t}\rVert^{2}}.$
@@ 4.3. User study setup
Figure 3. Screenshot from the data provided in the user study.
@@ 6. Correlation between objective metrics and subjective perception
Metric Kendall $\tau$ p-value 95% CI $\tau$
RMSE 0.2928 0.0267 (0.065, 0.334)
VD with GT 0.7345 3e-08 (0.618, 0.881)
VD 0.3684 0.0052 (0.149, 0.578)
BDP with GT 0.1538 0.2443 (-0.265, 0.283)
BDP 0.5464 3.358e-05 (0.44, 0.697)
Table 2. Kendall’s $\tau$ rank correlation coefficients and p-values for different objective metrics wrt the user study results. Bold font indicates statistically significant results (p¡0.01). We also include 95% confidence intervals for Kendall $\tau$ using non-parametric bootstrapping with 50,000 resamples, using the 2.5th and 97.5th percentile values.
@@
We also assess the inter-rater reliability using Krippendorff’s $\alpha$ (Krippendorff, 2018) with an ordinal distance function. This statistic was chosen because our rating scale consists of ordered categories (e.g., 1 = low, 5 = high), and $\alpha$ accounts for the magnitude of disagreement between raters—not just its occurrence. Importantly, Krippendorff’s $\alpha$ can accommodate unequal numbers of ratings per item, which applied in our case: not all stimuli were rated by the same number of participants. Confidence intervals were obtained via non-parametric bootstrapping (50,000 resamples), resampling at the stimulus level, and calculated using the 2.5th and 97.5th percentiles of the resulting $\alpha$ distribution. The resulting reliability estimate was $\alpha$ = 0.503 (95% CI: 0.346, 0.624). This suggests a moderate level of inter-rater reliability across study participants.